Every Monday, James has a math class and a biology class. The probability that he will skip his math class is 0.24, while the probability he will skip his biology class is 0.11. If the probability he will attend both classes is 0.69, what is the probability he will skip both classes?

Answer by ikleyn(52778)

.
The probability that he will attend Math class    = 1 - 0.24 = 0.76.


The probability that he will attend Biology class = 1 - 0.11 = 0.89.


The probability that he will attend both classes  = 0.69    ( ! given ! )


The probability that he will attend at least one of the two classes =

    0.76 + 0.89 - 0.69 = 0.96.    (*)


The probability that he will skip both classes is the complement to (*)

    1 - 0.96 = 0.04.


Answer.  The probability under the question is  0.04.

Solved.